# Towards Characterizing Subjectivity of Individuals through Modeling Value Conflicts and Trade-offs

### v1

@@ A.2 Clustering Value Conflicts and Trade-offs
We follow the approach used in Lam et al. (2024) to cluster the value representations.
@@
For the generated values that are conflicting in the situations in the dataset, we first obtain their vector representations. We use Open AI’s text embedding 3 small model (OpenAI, 2024), and reduced the dimensions to 256. We then further reduce the dimensionality using umap embeddings (McInnes et al., 2018). Figure 4 shows the 2-d visualization of umap embedding of all value representations, with number of neighbors as 30 and minimum distance as 0.
@@
After this step, we use HDBSCAN to generate initial clusters while enforcing minimum of 100 items in each cluster. After obtaining the initial cluster, we gather all uncategorized values. We then compute the distance between each of the uncategorized value and initial cluster representations. If the distance is closer than a threshold (0.95), we assign the item to the cluster. For values that are not close enough to any other clusters, we group them using Kmeans clustering.
@@
After assigning all value representations to a corresponding cluster, we ask gpt-4o-mini model to come up with a summary of unifying themes and patterns. Below is the template for the prompt:
@@
I have this set of bullet point summaries of text examples: {examples} Please write a summary of unifying patterns for these examples. For each high-level pattern, write a 5 word NAME for the pattern and an associated one-sentence ChatGPT PROMPT that could take in a new text example and determine whether the relevant pattern applies. Please also include 2 example_ids for items that BEST exemplify the pattern. Please respond ONLY with a valid JSON in the following format : {  "patterns": [   {    "name": ,    "prompt": ,    "example_ids": ,   }   {    "name": ,    "prompt": ,    "example_ids": ,   }  ] }
Appendix B Model Implementation
B.1 Hyperparameter Searching
For all fine-tuned language models, we perform hyperparameter searching on training batch size and learning rates, and the best working combination is 16 and 1.895e-5. When adopting LoRA, we also explored r and alpha values, and we find that $r=4$, $alpha=16$ works the best.
B.2 Computing Resources
For encoder-only models, fine-tuning 8 models for each redditor takes approximately 90 to 160 minutes on a single Tesla V100 GPU. Fine-tuning encoder-decoder models, BART and FLAN-T5, took around 4 hours and 6 hours, respectively.
@@ B.2 Computing Resources
For inference using gpt-4o models, each of the experiment costs around USD 20, where the costs for input query takes about USD 19.90 and the rest is for the output which is either 0 or 1.

### v2

@@ B.2 Clustering Value Conflicts and Trade-offs
The general framework of creating clusters from the annotated values follow the approaches suggested by Lam et al. (2024). The authors first cluster the texts using HDBSCAN, then expand the concepts by prompting LLMs. The initial clusters are formed using HDBSCAN (McInnes et al., 2017), and we later use LLMs to create additional clusters for values that remain uncategorized in the initial clustering phase, resulting in 111 clusters in total.
@@
For the generated values that are conflicting in the situations in the dataset, we first obtain their vector representations. We use OpenAI’s text-embedding-3-large model (OpenAI, 2024), and reduced the dimensions to 256. We then further reduce the dimensionality using umap embeddings (McInnes et al., 2018). Figure 6 shows the 2-D visualization of umap embedding of all value representations, with number of neighbors as 30 and minimum distance as 0.
@@
After this step, we use HDBSCAN to generate initial clusters while enforcing minimum of 100 items in each cluster. After obtaining the initial cluster, we gather all uncategorized values. We then compute the distance between each of the uncategorized value and initial cluster representations. If the distance is closer than a threshold (0.95), we assign the item to the cluster. For values that are not close enough to any other clusters, we group them using K-means clustering.
@@
After assigning all value representations to a corresponding cluster, we ask gpt-4o-mini model to come up with a summary of unifying themes and patterns. Below is the template for the prompt adopted from Lam et al. (2024):
Prompt for Expand Concepts I have this set of bullet point summaries of text examples:
{examples}
Please write a summary of unifying patterns for these examples.
@@
For each high-level pattern, write a 5 word NAME for the pattern and an associated one-sentence ChatGPT PROMPT that could take in a new text example and determine whether the relevant pattern applies.
Please also include 3 example_ids for items that BEST exemplify the pattern.
B.3 Comparison
Figure 7: Mapping results of LLM-generated values to Schwartz’s values. Numbers in each cell mean the log-sclaed count of the number of co-occurrences.
In order to analyze the usefulness of LLM-generated values and whether they cover the fixed set of values that are defined in a top-down approach, we map these values to the Schwartz’s values.
Figure 7 shows the mapping results. For each of the situation that has annotations of both LLM-generated values and Schwartz’s values, we count the co-occurrences between these values. For instance, if a situation’s annotated Schwartz’s value is “Security”, and at the same time, its annotated values that conflict is “Boundary Awareness and Respect”, then these two values have a co-occurrence. In the figure, we compute the log-scale of all the co-occurrence counts.
For each of the LLM-generated values, they cover multiple dimensions of Schwartz’s values, suggesting that these values have richer context information about the situations. Moreover, in some cases, these LLM-generated values cover conflicting Schwartz’s values together. For example, “Personal Boundaries and Well-being” value has high number of co-occurrences with “Security” and “Self-Direction”. These two Schwartz’s values conflict to each other, thus they are not categorized into the same bucket using the Schwartz’s values. This shows that LLM-generated values can be applied in a more flexible way, while preserving meaningful insights from the theories supported by Schwartz’s values.
Appendix C Model Implementation Details
C.1 Baseline Model Implementation
For all fine-tuned language models, we perform hyperparameter searching on training batch size and learning rates. For DistilBERT-base, the best working combination is 16 and 3.962e-5, for RoBERTa-base, it is 32 and 3.97838e-5, for DeBERTa-v3, it is 4 and 7.378218e-5. For encoder-only models, fine-tuning models for each redditor takes approximately 10 to 25 minutes on a single A30 GPU. Fine-tuning encoder-decoder models, BART and FLAN-T5, took around 15 to 30 minutes, respectively. Stretching this to 100 redditors and 5 fold experiments, running each model structure took a day to two days.
C.2 RAG-based Model Implementation
@@
For inference using GPT-4.1 models, each of the experiment costs around USD 30, where the costs for input query takes about USD 29 and the rest is for the output which is either 0 or 1. This was based on the batch prompting, which is half of the original price.
Appendix D Model Performance
Figure 8: Macro F1 score of fine-tuned models and RAG-based models. $x$-axis represents the attribute of redditors. The left figure shows how balanced the redditors’ judgment patterns are (more balanced if the value is closer to 1), and the right figure shows how many instances the redditors have.
Figure 8 illustrates the macro F1 score of the redditors varying in judgment distribution balance and the number of instances. For fine-tuned models, the F1 score decreases as the number of instances decreases and the judgment patterns become more skewed. For the RAG-based models, on the other hand, the F1 socre does not heavily depend on the judgment distributions and the number of instances for each redditor, as the RAG-based models can perform inference based on the few-shot examples only.